# Full-waveform Approximation of Finite-Sized Acoustic Apertures: Forward and Adjoint Wavefields

### v20

@@ 5.3. Reception Operator
We define the free-space operators $\mathcal{G}_{r}^{N}$ and $\mathcal{G}_{r}^{D}$, acting at an arbitrary but fixed time $t\in[0,T]$, on a field in free space as follows:
$\begin{split}\mathcal{G}_{r}^{N}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right)&\rightarrow C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right),\quad\mathcal{G}_{r}^{N}[f](\bm{x},t)=\rho_{0}\nabla\cdot\left(\frac{1}{\rho_{0}}\nabla f\right)(\bm{x},t),\\ \mathcal{G}_{r}^{D}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right)&\rightarrow C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right),\quad\mathcal{G}_{r}^{D}[f](\bm{x},t)=\frac{\partial f}{\partial\bm{n}_{r,+}}(\bm{x},t),\end{split}$ (56)
@@
$f:=\frac{1}{\rho_{0}}\frac{\partial p_{e}}{\partial\bm{n}},$
@@
$\lim_{b\to 0^{+}}\frac{1}{b}\,\Big[\frac{1}{\rho_{0}}\frac{\partial p_{e}}{\partial\bm{n}}\Big|_{\partial\nu_{r}^{+}}-\frac{1}{\rho_{0}}\frac{\partial p_{e}}{\partial\bm{n}}\Big|_{\partial\nu_{r}^{-}}\Big]=\nabla\cdot\Big(\frac{1}{\rho_{0}}\nabla p_{e}\Big),$ (59)
which, in the limit $b\to 0^{+}$, indicates the emergence of a distributional singularity in $\rho_{0}\,\nabla\cdot\!\big(\tfrac{1}{\rho_{0}}\nabla p_{e}\big)$ due to the sharp jump of $\tfrac{\partial p_{e}}{\partial\bm{n}_{r,+}}$ across the boundary, leading to the symmetric relation
$\lim_{b\to 0^{+}}\,\frac{2}{b}\,\tfrac{\partial p_{e}}{\partial\bm{n}_{r,+}}\Big|_{\partial\nu_{r}}\;=\;\rho_{0}\,\nabla\cdot\!\big(\tfrac{1}{\rho_{0}}\nabla p_{e}\big).$ (60)
This yields $y^{N}$ on the boundary as the action of the restriction operator on the div–grad field, as defined in Eq. (48) of Definition 9, where the surface-restricted Dirac delta distribution is given in Eq. (55).
@@
Remark 4.
In a practical setting and for the Dirichlet case, the pressure on the back surface of the transducer is controlled or set to zero, which leads to a singularity in the limit of very small thickness $h\to 0^{+}$ in the normal derivative of pressure across the thickness. This behaves like a spike, yielding
$\frac{\partial p_{e}}{\partial\bm{n}_{r,+}}\;\approx\;2\,\delta\!\left(0\right)\,p_{e}\Big|_{\partial\nu_{r}}\,.$ (61)
Consequently, replacing the Dirac delta with its regularized form yields Eq. (54), as established in Lemma 1. Since the thickness $h$ of the transducer is finite in practical settings, the normal pressure derivative is no longer a Dirac delta distribution, and the spread of the jump is controlled by $h$, rather than $b$. This leads to the restriction operator in Eq. (48) of Definition 9, where the surface-restricted Dirac delta distribution defined in Eq. 55 uses
@@
instead of $a_{R}=b$.
@@ 5.4. Forward Operator
where $\mathcal{P}$ is the wave propagation operator, and $\mathcal{R}_{r,\mathcal{G}}^{\{N,D\}}$ maps the resulting pressure wavefield in free space to the Neumann-type or Dirichlet-type boundary data, respectively. $y_{r,e}^{\{N,D\}}$ is then obtained as the action of $\mathcal{A}_{r}^{\{N,D\}}$ on the regularized source $\mathcal{S}_{e}$, which is induced by a boundary excitation on $\partial\nu_{e}\cap\partial\Omega$. Repeating this process for all emitter and receiver indices $e$ and $r$, respectively, yields the complete dataset.
@@ 6. Adjoint and Time-reversal Operators
As we analyze the integral formulations associated with finite-sized acoustic transducers, it is equally important to derive the corresponding adjoint operator under the same assumptions. This derivation relies on the integral representations introduced earlier.
@@ 6.1.2. Adjoint of the Reception Operator
Having derived the adjoint operator $\mathcal{P}^{*}$, we now proceed to derive the adjoint of the reception operators $\mathcal{R}_{r,\mathcal{G}}^{\{N,D\}}$, which are, respectively, composite operators involving the Neumann-type or Dirichlet-type free-space operators introduced in Definition 10, and the restriction operator introduced in Definition 9.
@@
$\begin{split}\left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}\!:&\;C_{0}^{\infty}\big(\partial\Omega\times[0,T]\big)\rightarrow C_{0}^{\infty}\big(\mathbb{R}^{d}\times[0,T]\big),\\ \left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}[f^{N}](\bm{x},t)&=\dfrac{a_{R}}{2}\displaystyle\int_{\partial\nu_{r}\cap\partial\Omega}dS(\bm{x}_{s})\,\nabla\cdot\left(\dfrac{1}{\rho_{0}}\nabla\left(\rho_{0}\delta_{b}(\bm{x}-\bm{x}_{s})\right)\right)f^{N}(\bm{x}_{s},t),\end{split}$ (69)
where $a_{R}$ is the bandwidth parameter of the regularized Dirac delta distribution, as defined below Eq. (55), which, in a practical setting, is replaced by the thickness of the active volume of the transducer, as defined in Eq. (62).
@@
At any fixed time $t\in[0,T]$, the adjoint operator $\left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}$, with respect to the standard $L^{2}$ bilinear form in $C_{0}^{\infty}(\partial\Omega\times[0,T])$ and $C_{0}^{\infty}(\mathbb{R}^{d}\times[0,T])$, must satisfy
$\begin{split}\frac{a_{R}}{2}\int_{\partial\nu_{r}\cap\partial\Omega}dS(\bm{x}_{s})\,f^{N}(\bm{x}_{s},t)\int_{\mathbb{R}^{d}}d\bm{x}\ \delta_{b}(\bm{x}_{s}-\bm{x})\,\rho_{0}(\bm{x})\nabla\cdot\left(\frac{1}{\rho_{0}(\bm{x})}\nabla f(\bm{x},t)\right)\\ =\int_{\mathbb{R}^{d}}d\bm{x}\ \left[\left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}[f^{N}](\bm{x}_{s},t)\right]f(\bm{x},t)\end{split}$ (70)
@@
$\int_{\mathbb{R}^{d}}d\bm{x}\left[\frac{a_{R}}{2}\int_{\partial\nu_{r}\cap\partial\Omega}dS(\bm{x}_{s})\,\nabla\cdot\left(\frac{1}{\rho_{0}}\nabla\left(\rho_{0}\delta_{b}(\bm{x}-\bm{x}_{s})\right)\right)f^{N}(\bm{x}_{s},t)\right]f(\bm{x},t).$ (71)
By comparing Eq. (71) with the right-hand side of Eq. (70), we identify the bracketed term as $\left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}[f^{N}]$, thereby confirming the adjoint expression in Eq. (69). ∎
@@
At any fixed time $t\in[0,T]$, the adjoint operator $\left(\mathcal{R}_{\mathcal{G},r}^{D}\right)^{*}$, with respect to the standard $L^{2}$ bilinear form in $C_{0}^{\infty}(\partial\Omega\times[0,T])$ and $C_{0}^{\infty}(\mathbb{R}^{d}\times[0,T])$, must satisfy
$\begin{split}\frac{a_{R}}{2}\int_{\partial\nu_{r}\cap\partial\Omega}dS(\bm{x}_{s})\,f^{D}(\bm{x}_{s},t)\int_{\mathbb{R}^{d}}d\bm{x}\ \delta_{b}(\bm{x}_{s}-\bm{x})\ \big[\nabla f(\bm{x},t)\cdot\bm{n}_{r,+}\big],\\ =\int_{\mathbb{R}^{d}}d\bm{x}\ \left[\left(\mathcal{R}_{\mathcal{G},r}^{D}\right)^{*}[f^{D}](\bm{x}_{s},t)\right]f(\bm{x},t)\end{split}$ (73)
@@
By comparing Eq. (74) with the right-hand side of Eq. (73), we identify the bracketed term as $\left(\mathcal{R}_{\mathcal{G},r}^{D}\right)^{*}[f^{D}]$, thereby confirming the adjoint expression in Eq. (72). ∎
@@
The action of the adjoint of the forward operator $\mathcal{A}_{r}$, introduced in Definition 13, on any measured boundary data $y_{r,e}^{\{N,D\}}\in C_{0}^{\infty}\big(\partial\Omega\times[0,T]\big)$ is given by
$\mathcal{A}_{r}^{*}\left[y_{r,e}^{\{N,D\}}\right](\bm{x},t)=p_{e,r}^{*}(\bm{x}^{\prime},T-t^{\prime}),$ (75)
@@
$p_{e,r}^{*}(\bm{x}^{\prime},0)=0,\quad\frac{\partial}{\partial t^{\prime}}p_{e,r}^{*}(\bm{x}^{\prime},0)=0,\quad\text{in}\quad\mathbb{R}^{d}.$ (77)
@@
The adjoint operator $\mathcal{A}_{r}^{*}$ is composed of the adjoints of the constituent operators:
$\mathcal{A}_{r}^{*}=\mathcal{P}^{*}\left(\mathcal{R}_{\mathcal{G},r}^{\{N,D\}}\right)^{*},$ (79)

### v21

@@ 5.3. Reception Operator
We define the free-space operators $\mathcal{G}_{r}^{N}$ and $\mathcal{G}_{r}^{D}$, acting on a field in free space as:
$\begin{split}\mathcal{G}_{r}^{N}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right)&\rightarrow C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right),\quad\mathcal{G}_{r}^{N}[f](\bm{x},t):=\rho_{0}\nabla\cdot\left(\frac{1}{\rho_{0}}\nabla f(\bm{x},t)\right),\\ \mathcal{G}_{r}^{D}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right)&\rightarrow C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right),\quad\mathcal{G}_{r}^{D}[f](\bm{x},t):=\frac{\partial f}{\partial\bm{n}_{r,+}}(\bm{x},t),\end{split}$ (59)
@@
$f:=\frac{1}{\rho_{0}}\frac{\partial p_{e}}{\partial\bm{n}_{r,+}},$
@@
$\lim_{b\to 0^{+}}\frac{1}{b}\,\Big[\frac{1}{\rho_{0}}\frac{\partial p_{e}}{\partial\bm{n}_{r,+}}\Big|_{\partial\nu_{r}^{+}}-\frac{1}{\rho_{0}}\frac{\partial p_{e}}{\partial\bm{n}_{r,+}}\Big|_{\partial\nu_{r}^{-}}\Big](\bm{x}_{s},t)\;=\;\int_{\mathbb{R}^{d}}d\bm{x}\ \delta_{b}(\bm{x}_{s}-\bm{x})\nabla\cdot\Big(\frac{1}{\rho_{0}}\nabla p_{e}(\bm{x},t)\Big),$ (62)
which, in the limit $b\to 0^{+}$, indicates the emergence of a distributional singularity in $\rho_{0}\,\nabla\cdot\!\big(\tfrac{1}{\rho_{0}}\nabla p_{e}\big)$ due to the sharp jump of $2\tfrac{\partial p_{e}}{\partial\bm{n}_{r,+}}$ across the boundary, leading to the symmetric relation
$\lim_{b\to 0^{+}}\,\frac{2}{b}\,\tfrac{\partial p_{e}}{\partial\bm{n}_{r,+}}\Big|_{\partial\nu_{r}}(\bm{x}_{s},t)\;=\;\int_{\mathbb{R}^{d}}d\bm{x}\ \delta_{b}(\bm{x}_{s}-\bm{x})\rho_{0}\nabla\cdot\Big(\frac{1}{\rho_{0}}\nabla p_{e}(\bm{x},t)\Big).$ (63)
This yields $y^{N}$ on the boundary as the action of the restriction operator, defined in Eq. (51) of Definition 9, applied to the div-grad field, where the surface-restricted Dirac delta distribution is given in Eq. (58).
@@
Reception Operator in Practical Setting
In practical settings, and for the Dirichlet case, the pressure within the transducer crystal decays sharply, approaching zero at the backing surface. This produces a smeared singularity in the normal derivative of the pressure across the thickness as $h\to 0^{+}$. In the limit of very small but finite thickness, this behaves like a spike, yielding
$2\,p_{e}\Big|_{\text{crystal}}=\int_{-\frac{h}{2}}^{+\frac{h}{2}}\frac{\partial p_{e}}{\partial\bm{n}_{r,+}}\,d\zeta\longleftrightarrow 2\,p_{e}\Big|_{\partial\nu_{r}}(\bm{x}_{s},t)\;\approx\;h\int_{\mathbb{R}^{d}}d\bm{x}\,\delta_{b}(\bm{x}_{s}-\bm{x})\frac{\partial p_{e}}{\partial\bm{n}_{r,+}}(\bm{x},t),$ (64)
where $\zeta$ denotes the normal coordinate. For the right-hand-side integral expression, we assume that the normal derivative of the pressure field remains approximately constant across the small but finite thickness of the thin layer. Recall that, although the volume integral is formally taken over free space, its effective contribution is confined to a thin layer surrounding the infinite plane—corresponding, in practice, to the finite thickness of the crystal.
Comparing the right-hand-side integral expression in (64) with (57), we note that, since the crystal thickness $h$ is finite in practice, the normal contribution of the surface-restricted Dirac delta distribution does not exhibit a jump controlled by $b$; instead, the spread of the jump is determined by $h$. This motivates the restriction operator in Eq. (51) of Definition 9, where the surface-restricted Dirac delta distribution defined in Eq. (58) is generalized to
@@
with the special case $a_{R}=b$ in the ideal limit $h\to 0^{+}$.
@@ 5.4. Forward Operator
where $\mathcal{P}$ is the wave propagation operator, and $\mathcal{R}_{r,\mathcal{G}}^{\{N,D\}}$ maps the resulting pressure wavefield in free space to the Neumann-type or Dirichlet-type boundary data, respectively. $y_{r,e}^{\{N,D\}}$ then corresponds to the action of $\mathcal{A}_{r}^{\{N,D\}}$ on the regularized source $\mathcal{S}_{e}$, which is induced by a boundary excitation on $\partial\nu_{e}\cap\partial\Omega$. Repeating this process for all emitter and receiver indices $e$ and $r$, respectively, yields the complete dataset.
@@
Recall from Section 5.1.2 that, in the particular case of photoacoustic tomography, the operator $\mathcal{A}_{r}$ acts on a source $s_{\mathcal{Q}}$, which is itself determined by the initial pressure distribution $p_{0}$ through the mapping defined in Eq. (40).
@@ 6. Adjoint and Time-reversal Operators
In conjunction with the integral formulations governing the emission and reception processes of finite-sized acoustic transducers, it is essential to derive the corresponding adjoint operator under the same assumptions. The resulting formulation follows directly from the integral representations established in the preceding sections.
@@ 6.1.2. Adjoint of the Reception Operator
Having derived the adjoint operator $\mathcal{P}^{*}$, we now proceed to derive the adjoint of the reception operators $\mathcal{R}_{r,\mathcal{G}}^{\{N,D\}}$, which are, respectively, composite operators consisting of the Neumann-type or Dirichlet-type free-space operators introduced in Definition 10, and the restriction operator introduced in Definition 9.
@@
$\begin{split}\left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}\!:&\;C_{0}^{\infty}\big(\partial\Omega\times[0,T]\big)\rightarrow C_{0}^{\infty}\big(\mathbb{R}^{d}\times[0,T]\big),\\ \left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}[f^{N}](\bm{x},t)&=\dfrac{a_{R}}{2}\displaystyle\int_{\partial\nu_{r}\cap\partial\Omega}dS(\bm{x}_{s})\,\nabla\cdot\left(\dfrac{1}{\rho_{0}}\nabla\Big(\rho_{0}\delta_{b}(\bm{x}-\bm{x}_{s})\Big)\right)f^{N}(\bm{x}_{s},t),\end{split}$ (72)
where $a_{R}$ denotes the thickness $h$ of the active volume (crystal) of the receiver, as defined in Eq. (65). In the idealized case $h\to 0^{+}$, $a_{R}$ represents instead the bandwidth parameter of the regularized Dirac delta distribution, as defined in Eq. (58).
@@
At any arbitrary but fixed time $t\in[0,T]$, the adjoint operator $\left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}$, with respect to the standard $L^{2}$ bilinear form in $C_{0}^{\infty}(\partial\Omega\times[0,T])$ and $C_{0}^{\infty}(\mathbb{R}^{d}\times[0,T])$, must satisfy
$\begin{split}\frac{a_{R}}{2}\int_{\partial\nu_{r}\cap\partial\Omega}dS(\bm{x}_{s})\,f^{N}(\bm{x}_{s},t)\int_{\mathbb{R}^{d}}d\bm{x}\ \delta_{b}(\bm{x}_{s}-\bm{x})\,\rho_{0}(\bm{x})\nabla\cdot\left(\frac{1}{\rho_{0}(\bm{x})}\nabla f(\bm{x},t)\right)\\ =\int_{\mathbb{R}^{d}}d\bm{x}\ \left[\left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}[f^{N}](\bm{x},t)\right]f(\bm{x},t)\end{split}$ (73)
@@
$\int_{\mathbb{R}^{d}}d\bm{x}\left[\frac{a_{R}}{2}\int_{\partial\nu_{r}\cap\partial\Omega}dS(\bm{x}_{s})\,\nabla\cdot\left(\frac{1}{\rho_{0}}\nabla\Big(\rho_{0}\delta_{b}(\bm{x}-\bm{x}_{s})\Big)\right)f^{N}(\bm{x}_{s},t)\right]f(\bm{x},t).$ (74)
By comparing the integral expression in (74) with the right-hand side of Eq. (73), we identify the bracketed term as $\left(\mathcal{R}_{\mathcal{G},r}^{N}\right)^{*}[f^{N}]$, thereby confirming the adjoint expression in Eq. (72). ∎
@@
At any arbitrary but fixed time $t\in[0,T]$, the adjoint operator $\left(\mathcal{R}_{\mathcal{G},r}^{D}\right)^{*}$, with respect to the standard $L^{2}$ bilinear form in $C_{0}^{\infty}(\partial\Omega\times[0,T])$ and $C_{0}^{\infty}(\mathbb{R}^{d}\times[0,T])$, must satisfy
$\begin{split}\frac{a_{R}}{2}\int_{\partial\nu_{r}\cap\partial\Omega}dS(\bm{x}_{s})\,f^{D}(\bm{x}_{s},t)\int_{\mathbb{R}^{d}}d\bm{x}\ \delta_{b}(\bm{x}_{s}-\bm{x})\ \big[\nabla f(\bm{x},t)\cdot\bm{n}_{r,+}\big],\\ =\int_{\mathbb{R}^{d}}d\bm{x}\ \left[\left(\mathcal{R}_{\mathcal{G},r}^{D}\right)^{*}[f^{D}](\bm{x},t)\right]f(\bm{x},t)\end{split}$ (76)
@@
By comparing the integral expression in (77) with the right-hand side of Eq. (76), we identify the bracketed term as $\left(\mathcal{R}_{\mathcal{G},r}^{D}\right)^{*}[f^{D}]$, thereby confirming the adjoint expression in Eq. (75). ∎
@@
The action of the adjoint of the forward operator $\mathcal{A}_{r}^{\{N,D\}}$, introduced in Definition 13, on any measured boundary data $y_{r,e}^{\{N,D\}}\in C_{0}^{\infty}\big(\partial\Omega\times[0,T]\big)$ is given by
$\left(\mathcal{A}_{r}^{\{N,D\}}\right)^{*}\left[y_{r,e}^{\{N,D\}}\right](\bm{x}^{\prime},t^{\prime})=p_{e,r}^{*}(\bm{x}^{\prime},T-t^{\prime}),$ (78)
@@
$p_{e,r}^{*}(\bm{x}^{\prime},0)=0,\quad\frac{\partial}{\partial t^{\prime}}p_{e,r}^{*}(\bm{x}^{\prime},0)=0,\quad\bm{x}^{\prime}\in\mathbb{R}^{d}.$ (80)
@@
The adjoint operator $\left(\mathcal{A}_{r}^{\{N,D\}}\right)^{*}$ is composed of the adjoints of the constituent operators:
$\left(\mathcal{A}_{r}^{\{N,D\}}\right)^{*}=\mathcal{P}^{*}\left(\mathcal{R}_{\mathcal{G},r}^{\{N,D\}}\right)^{*},$ (82)
@@
6.1.3. Adjoint Operator in Photoacoustic Tomography
Recall from Section 5.1.2 that the photoacoustic forward operator that maps to the Dirichlet boundary data is defined as the composite map $\mathcal{A}_{r}^{D}\circ\mathcal{Q}$, acting on the initial pressure distribution $p_{0}\in C_{0}^{\infty}(\Omega^{-})$ to produce the measured Dirichlet boundary data, i.e., $y_{r}^{D}\in C_{0}^{\infty}(\partial\Omega\times[0,T])$. (Since $p_{0}$ is independent of the index $e$, this dependence is omitted throughout the formulation.) Correspondingly, the adjoint of the photoacoustic forward operator is given by the composite map $\mathcal{Q}^{*}\circ\left(\mathcal{A}_{r}^{D}\right)^{*}$, acting on the Dirichlet boundary data $y_{r}^{D}$.
Lemma 8.
The action of the adjoint of the operator $\mathcal{Q}$ on any test function $f^{s_{\mathcal{Q}}}\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{+})$ is given by
$\begin{split}&\mathcal{Q}^{*}:C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{+})\rightarrow C_{0}^{\infty}(\Omega^{-}),\\ &\mathcal{Q}^{*}[f^{s_{\mathcal{Q}}}]=-\frac{1}{c(\bm{x})^{2}}\,\frac{\partial f^{s_{\mathcal{Q}}}}{\partial t}(\bm{x},0).\end{split}$ (83)
Proof.
The adjoint operator $\mathcal{Q}^{*}$ with respect to the standard $L^{2}$ bilinear form in $C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{+})$ and $C_{0}^{\infty}(\Omega^{-})$ is defined by
$\int_{\mathbb{R}^{d}}d\bm{x}\int_{\mathbb{R}^{+}}dt\,f^{s_{\mathcal{Q}}}(\bm{x},t)\,\frac{1}{c(\bm{x})^{2}}f(\bm{x})\,\frac{d}{dt}\delta(t)=\int_{\Omega^{-}}d\bm{x}\,\Big[\mathcal{Q}^{*}[f^{s_{\mathcal{Q}}}](\bm{x})\Big]\,f(\bm{x})$ (84)
for any $f^{s_{\mathcal{Q}}}\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{+})$ and $f\in C_{0}^{\infty}(\Omega^{-})$.
Applying integration by parts in time to the left-hand side of (84) yields
$\int_{\Omega^{-}}d\bm{x}\,\Big[-\frac{1}{c(\bm{x})^{2}}\frac{\partial f^{s_{\mathcal{Q}}}}{\partial t}(\bm{x},0)\Big]f(\bm{x}),$ (85)
where we have also restricted the domain of integration to the support of $f$.
Comparing the integral expression in (85) with the right-hand side of Eq. (84) immediately identifies the bracketed term in (85) as $\mathcal{Q}^{*}[f^{s_{\mathcal{Q}}}]$, which proves the expression in (83). ∎
Consequently, the full composite adjoint operator for photoacoustic tomography is given by
$\left(\mathcal{A}_{r}^{D}\right)^{*}[y_{r}^{D}](\bm{x}^{\prime})=-\frac{1}{c(\bm{x})^{2}}\frac{\partial p_{r}^{*}}{\partial t}(\bm{x}^{\prime},T),$ (86)
where $p_{r}^{*}$ satisfies the adjoint wave equation (79), with the Cauchy conditions (80) and with a time-reversed adjoint source $\left(\mathcal{S}^{*}\right)_{r}^{D}$ that satisfies the second line in Eq. (81). (Recall that the dependence on $e$ should be omitted throughout these formulas.)